# Discrete Diffusion Modeling by Estimating the Ratios of the Data Distribution

### v2

@@ Abstract
Despite their groundbreaking performance for many generative modeling tasks, diffusion models have fallen short on discrete data domains such as natural language. Crucially, standard diffusion models rely on the well-established theory of score matching, but efforts to generalize this to discrete structures have not yielded the same empirical gains. In this work, we bridge this gap by proposing score entropy, a novel loss that naturally extends score matching to discrete spaces, integrates seamlessly to build discrete diffusion models, and significantly boosts performance. Experimentally, we test our Score Entropy Discrete Diffusion models (SEDD) on standard language modeling tasks. For comparable model sizes, SEDD beats existing language diffusion paradigms (reducing perplexity by -%) and is competitive with autoregressive models, in particular outperforming GPT-2. Furthermore, compared to autoregressive mdoels, SEDD generates faithful text without requiring distribution annealing techniques like temperature scaling (around - better generative perplexity than un-annealed GPT-2), can trade compute and quality (similar quality with fewer network evaluations), and enables controllable infilling (matching nucleus sampling quality while enabling other strategies besides left to right prompting).†† We open source our code at github.com/louaaron/Score-Entropy-Discrete-Diffusion
@@ 1 Introduction
To alleviate these issues, researchers have sought alternative approaches to generating text data. In particular, inspired by their success in the image domain, many works have extended diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b) to language domains (Li et al., 2022; Austin et al., 2021). Yet, despite considerable effort, no such approach yet rivals autoregressive modeling, as they are not competitive on likelihoods, are slower to sample from, and do not generate comparable samples without resorting to heavy annealing and empirical alterations.
In our work, we challenge the longstanding dominance of autoregressive models by introducing Score Entropy Discrete Diffusion models (SEDD). SEDD parameterizes a reverse discrete diffusion process using the ratios of the data distribution. These are learned using score entropy, a novel loss that is analogous to score matching for standard diffusion models (Hyvärinen, 2005; Song & Ermon, 2019) and results in several empirical benefits:
@@
On core language modeling tasks, SEDD outperforms existing all language diffusion models (Li et al., 2022; Austin et al., 2021; Gulrajani & Hashimoto, 2023; He et al., 2022) by large margins and is competitive with autoregressive models of the same size (beating GPT-2 on its zero-shot perplexity tasks (Radford et al., 2019)).
@@ 3 Score Entropy Discrete Diffusion Models
In this section, we introduce score entropy. Similar to concrete score matching, we learn the collected concrete score (—). We design the score entropy loss to incorporate the fact that these ratios are positive and evolve under a discrete diffusion.
@@ 4.2 Arbitrary Prompting and Infilling
where is concatenation along and . Since the unconditional and conditional scores coincide, we can use our (learned unconditionally) for conditional sampling (given arbitrary ). For a -leaping update rule (Equation 26 or 28), one would only modify by changing the values at . An explicit pseudocode of this is given in Algorithm 3.
@@ 5.2.2 One Billion Words Dataset
| Ours (NAR) | SEDD Uniform | 1.41 |
| SEDD Absorb | 1.32 |
@@ 5.3.1 Unconditional Generation
For SEDD, we simulate using 32 to 2048 steps, which approximates the learned distribution with minimal error for a large number of steps (the sequences are length 1024). Our results (both the measured generative perplexity and some samples) are shown in Figure 1. SEDD matches GPT-2 quality using 32 fewer network evaluations and outperforms by - when using the full 2048 steps. Furthermore, SEDD forms a predictable log-log linear pareto frontier between the number of sampling steps and generative perplexity. While we did not run a systemic evaluation of runtime (as this would be out of scope), SEDD with 100 steps takes about as long as GPT-2 with KV-caching on an A100 GPU.
@@ 6 Related Work
Discrete Diffusion Models. Most discrete diffusion works follow the framework set out by D3PM (Austin et al., 2021) which mimics “mean prediction” (Ho et al., 2020). However, these are largely applied to fields other than language (e.g. images), likely due to empirical challenges. Despite this, some works have shown strong performance on language, particularly for seq-to-seq tasks and more efficient generation (Zheng et al., 2023; Chen et al., 2023; Ye et al., 2023). In particular, discrete diffusion is promising because, unlike continuous language diffusion, there is ample evidence suggesting that one can sample with fewer steps.
SEDD vs Prior Work. SEDD is a discrete diffusion model that focuses on score matching, the crucial ingredient for continuous diffusions (Song & Ermon, 2019; Ho et al., 2020). While there are some conceptual similarities with some prior work (Campbell et al., 2022; Sun et al., 2023) (since all methods seek to reverse the same discrete diffusion processes), SEDD focuses on a principled, scalable, and performant objective, filling in corresponding shortcomings found in previous works. These critical differences enable large improvements for language tasks, where prior discrete diffusion models have conspicuously struggled on.
@@ 7 Conclusion
8 Acknowledgements
@@ Appendix A Proof of Main Results
Effectively, is the path measure KL divergence (Campbell et al., 2022; Song et al., 2021a), and the proof follows similarly. In particular, we have that, by the data processing inequality

### v3

@@ Abstract
Despite their groundbreaking performance for many generative modeling tasks, diffusion models have fallen short on discrete data domains such as natural language. Crucially, standard diffusion models rely on the well-established theory of score matching, but efforts to generalize this to discrete structures have not yielded the same empirical gains. In this work, we bridge this gap by proposing score entropy, a novel loss that naturally extends score matching to discrete spaces, integrates seamlessly to build discrete diffusion models, and significantly boosts performance. Experimentally, we test our Score Entropy Discrete Diffusion models (SEDD) on standard language modeling tasks. For comparable model sizes, SEDD beats existing language diffusion paradigms (reducing perplexity by -%) and is competitive with autoregressive models, in particular outperforming GPT-2. Furthermore, compared to autoregressive mdoels, SEDD generates faithful text without requiring distribution annealing techniques like temperature scaling (around - better generative perplexity than un-annealed GPT-2), can trade compute and quality (similar quality with fewer network evaluations), and enables controllable infilling (matching nucleus sampling quality while enabling other strategies besides left to right prompting).
@@ 1 Introduction
To alleviate these issues, researchers have sought alternative approaches to generating text data. In particular, inspired by their success in the image domain, many works have extended diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021c) to language domains (Li et al., 2022; Austin et al., 2021). Yet, despite considerable effort, no such approach yet rivals autoregressive modeling, as they are not competitive on likelihoods, are slower to sample from, and do not generate comparable samples without resorting to heavy annealing and empirical alterations.
In our work, we challenge the longstanding dominance of autoregressive models by introducing Score Entropy Discrete Diffusion models (SEDD). SEDD parameterizes a reverse discrete diffusion process using the ratios of the data distribution. These are learned using score entropy, a novel loss that is analogous to score matching for standard diffusion models (Hyvärinen, 2005; Song & Ermon, 2019) and results in several empirical benefits††We open source our code at github.com/louaaron/Score-Entropy-Discrete-Diffusion:
@@
On core language modeling tasks, SEDD outperforms all existing language diffusion models (Li et al., 2022; Austin et al., 2021; Gulrajani & Hashimoto, 2023; He et al., 2022) by large margins and is competitive with autoregressive models of the same size (beating GPT-2 on its zero-shot perplexity tasks (Radford et al., 2019)).
@@ 3 Score Entropy Discrete Diffusion Models
In this section, we introduce score entropy. Similar to concrete score matching, we learn the collected concrete score (). We design the score entropy loss to incorporate the fact that these ratios are positive and evolve under a discrete diffusion.
@@ 4.2 Arbitrary Prompting and Infilling
where is concatenation along and . Since the unconditional and conditional scores coincide, we can use our (learned unconditionally) for conditional sampling (given arbitrary ). For a -leaping update rule (Equation 17 or 19), one would only modify by changing the values at . An explicit pseudocode of this is given in Algorithm 3.
@@ 5.2.2 One Billion Words Dataset
| Ours (NAR) | SEDD Uniform | 1.47 |
| SEDD Absorb | 1.39 |
@@ 5.3.1 Unconditional Generation
For SEDD, we simulate using 32 to 2048 steps, which approximates the learned distribution with minimal error for a large number of steps (the sequences are length 1024). Our results (both the measured generative perplexity and some samples) are shown in Figure 1. SEDD matches GPT-2 quality using 32 fewer network evaluations and outperforms by - when using the full 2048 steps. Furthermore, SEDD forms a predictable log-log linear pareto frontier between the number of sampling steps and generative perplexity. However, each network evaluation is different due to the KV-cache, which introduces a cost benefit tradeoff that we discuss more in Section 6.
@@ 6 Related Work
Discrete Diffusion Models. Most discrete diffusion works follow the framework set out by D3PM (Austin et al., 2021) which mimics “mean prediction” (Ho et al., 2020). These discrete diffusion methods are largely applied to fields other than language (e.g. images), likely due to empirical challenges. Despite this, some works have shown strong performance on language, particularly for seq-to-seq tasks and more efficient generation (Zheng et al., 2023; Chen et al., 2023; Ye et al., 2023). Notably, from these works discrete diffusion has tended to be advantageous over continuous diffusion in reducing network evaluations.
SEDD vs Prior Work. SEDD is a discrete diffusion model that focuses on score matching, the crucial ingredient for continuous diffusions (Song & Ermon, 2019; Ho et al., 2020). Many such works also focus on reversing a discrete diffusion process (Campbell et al., 2022; Benton et al., 2022; Sun et al., 2023), so score entropy is naturally related with prior training objectives. However, SEDD focuses on a principled, scalable, and performant objective (namely denoising score entropy), filling in shortcomings found in previous works. In particular, prior methods train either with the equivalent of implicit score entropy (which is intractable and high variance) or propose alternate losses that suffer from other issues. These critical differences enable large improvements for language tasks, where prior discrete diffusion models have conspicuously struggled on.
@@
SEDD vs Autoregressive Sampling Iterations. SEDD and autoregressive models have significantly different sampling procedures due to the introduction of the KV-cache for standard decoder-only transformer models. In particular, this complicates the inference code (as each network pass changes from being a standard full batch forward) and trades off speed with memory. For example, for our (known) unoptimized codebase and the existing huggingface transformers library (Wolf et al., 2020), we observed that SEDD matches autoregressive inference time when using around 100 steps but can increase the batch size by roughly times by removing the KV-cache memory. Future work will likely decrease the steps required for optimal generation (similar to existing work in standard diffusion (Song et al., 2021a)) which can improve this tradeoff.
@@ 7 Conclusion
Impact Statement
This paper proposes work that advances the field of natural language generation. Outside of existing ethical questions for this area (e.g. bias, toxicity, fake content), our approach does not present any specific danger as the core work is largely theoretical and not at the scale to pose a specific problem.
Acknowledgements
@@ Appendix A Proof of Main Results
Effectively, is the path measure KL divergence (Campbell et al., 2022; Song et al., 2021b), and the proof follows similarly. In particular, we have that, by the data processing inequality